If a piece of real estate purchased for $50,000 in 1998 appreciates at the rate of 5% per year, then its value (t) years after the purchase will be f(t)=50,000(1.05^t). According to this model, by how much will the value of this piece of property increase between the years 2007 and 2008?
Answer by Fermat(127)

f(t) = 50,000(1.05)^t
Property is bought in 1998.
By the beginning of 2007, 9 years have passed.
By the beginning of 2008, 10 years have passed.
Value of property by beginning of 2007 is f(9).
f(9) = 50,000(1.05)^9 = 77,566.41
Value of property by beginning of 2008 is f(10).
f(10) = 50,000(1.05)^10 = 81,444.73
Increase in value during 2007 is: 81,444.73 - 77,566.41 = 3,878.32
Ans: $3,878.32